OpenStudy (anonymous):

Express the complex number in trigonometric form. -3i?

OpenStudy (psymon):

Guess it wants polar form, lol. Well, rectangular form for a complex number is: a + bi the polar form (trig form) is: \[r(\cos \theta + isin \theta)\] So now there are various conversions we have to go back and forth between the two. We need two of them, but I'll list all 4 since you'll need the others for different problems

OpenStudy (psymon):

Rectangular to polar: \[r= \sqrt{a ^{2}+b ^{2}}\] \[\tan \theta = \frac{ b }{ a }\] Polar to rectangular: \[x = rcos \theta \] \[y = rsin \theta \] So we'll use the top two conversions and go from there. With me so far?

OpenStudy (anonymous):

yes!!

OpenStudy (psymon):

Alright, so what is a and what is b in your problem?

OpenStudy (anonymous):

it only says for -3i

OpenStudy (psymon):

Right. So if you have the form of a + bi, but you only have bi, what must a be?

OpenStudy (anonymous):

0!

OpenStudy (psymon):

Bingo :P

OpenStudy (psymon):

So now we need theta and r so we can put it in polar form. So we must do the 1st conversion to get r. \[r = \sqrt{0^{2} + (-3)^{2}} = ?\] So that gives me?

OpenStudy (anonymous):

3 haha

OpenStudy (psymon):

haha, right. So r is 3. Now we need theta, which weget using: \[\tan \theta = \frac{ b }{ a }\]Problem is we justsaid a was 0, so we have: \[\tan \theta = \frac{ -3 }{ 0 }\], which is undefined. So basically, where is tangent(theta) undefined?

OpenStudy (anonymous):

so is it 0?

OpenStudy (psymon):

No, tan0 is 0, its defined there.

OpenStudy (anonymous):

i get everything up to this point :/

OpenStudy (psymon):

Well: \[\tan \theta = \frac{ \sin \theta }{ \cos \theta } \] So if we want this to be undefined, we want values where cos(theta) = 0, that way we get 0 in the denominator.

OpenStudy (anonymous):

cos = 0 at 270

OpenStudy (anonymous):

and 90

OpenStudy (psymon):

Yep, its 0 at 90 and 270. But we know it's 270 because we have 0 -3i. 0 - 3i is the same as (0,-3), which means we are at y = -3, which definitely means 270. So 270 or 3pi/2 is your theta, giving us: \[3(\cos270 + isin270)\] And that's your answer :3

OpenStudy (anonymous):

ahh thank you! do you think you can help me with three more problems? they're all like this i just wanna make sure i'm right

OpenStudy (psymon):

Lol, alright then xD

OpenStudy (anonymous):

Express the complex number in trigonometric form. -3 +3rad3 i

OpenStudy (anonymous):

so i did the first part and got 3rad3

OpenStudy (psymon):

For your r?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

what do i do from there?

OpenStudy (psymon):

Sorry. And you should get 6 for the radius xD

OpenStudy (anonymous):

ugh typical i messed that up okay an then what do i do again? :(

OpenStudy (psymon):

Alright, so r = 6. Now we need to do: \[\tan \theta = \frac{ b }{ a }\] So foryou, you have: \[\tan \theta= \frac{ 3\sqrt{3} }{ -3 }\] So now you just need to find the angle of tangent that equals that.

OpenStudy (anonymous):

how do i figure that out

OpenStudy (psymon):

Well, it simplifies to -sqrt(3), right?

OpenStudy (anonymous):

oh crap yes

OpenStudy (psymon):

Mhm. Well that's actually a clean value for tangent xD

OpenStudy (anonymous):

so would it be 5pi/6

OpenStudy (anonymous):

answers are in radians

OpenStudy (psymon):

Not quite. Correct quadrant, though :P

OpenStudy (anonymous):

other option is 2pi/3 lmao

OpenStudy (psymon):

Haha, well if you have multiple choice then that makes it easy, eh?

OpenStudy (anonymous):

yep!! okay two more than i promise i will let you be haha

OpenStudy (psymon):

Fair enough xD

OpenStudy (anonymous):

Write the complex number in the form a + bi. 5/2(cos 150° + i sin 150°)

OpenStudy (psymon):

These are actually easier. What is the cosine and sine of 150?

OpenStudy (anonymous):

-rad3/2 and 1/2

OpenStudy (psymon):

Okay, so now multiply both of those by (5/2)

OpenStudy (anonymous):

-5rad3/4, 5/4 oh wow that's so easy

OpenStudy (psymon):

Just make sure you have an "i" on the 5/4 part. That i doesnt disappear xD

OpenStudy (anonymous):

it's multiple choice :) okay last one is different so if you dont know how to do it its alll good

OpenStudy (psymon):

Alright, haha. I hope I do xD

OpenStudy (anonymous):

i'm sure you know though

OpenStudy (psymon):

never though

OpenStudy (anonymous):

Two forces with magnitudes of 25 and 30 pounds act on an object at angles of 10° and 100° respectively. Find the direction and magnitude of the resultant force. Round to two decimal places in all intermediate steps and in your final answer.

OpenStudy (psymon):

never know*

OpenStudy (psymon):

Tricky :P But yeah, might be better if I draw this.

OpenStudy (anonymous):

yeah i had another one like this but it was waayy easier

OpenStudy (psymon):

Hmm....nvm, no drawing, not sure how to draw it, haha.

OpenStudy (anonymous):

haha if u can't get it it's alright

OpenStudy (psymon):

Okay, so basically, if you push on an object directly with 25 pounds of force, the object will feel the full 25 pounds of force. But if you push with 25 pounds of force at an angle, the object only receives a fraction of that force. You put more energy in than the object takes. So depending on the angle, the amount of power is reduced. In order to find out how much of a force is actually being felt, we need the polar equations we've been using. So for example, we have a 25 pound force at a 10 degree angle. This means the force that the object actually feels is 25(cos(10) + sin(10))

OpenStudy (psymon):

So what I am going to do is mark cosine with the variable i and sin with the variable j (because usually i is x-direction and j is y-direction with vectors). So the way to solve this is to add the two equations. The first equation I just gave: \[25(\cos(10)i + \sin(10)j)\] The 2nd force equation is \[30(\cos(100)i + \sin(100)j)\] Now don't do the calculator work or anything, just add those two equations together \[25(\cos(10)i+\sin(10)j) + 30(\cos(100)i + \sin(100)j)\]

OpenStudy (psymon):

Well, guess there is no way to add them, the angles aren't the same. Lol, just multiply it out and such, haha.

OpenStudy (psymon):

And don't drop the i and j o.o

OpenStudy (anonymous):

multiply them or add?

OpenStudy (psymon):

Well, distribute the 25 and the 30. Then it's all calculator work. You'll have 2 things that have the variable i and two things with the variable j that you'll combine because of like terms.

OpenStudy (anonymous):

is it 704

OpenStudy (psymon):

What happened to i and j? xD

OpenStudy (anonymous):

hahaha ahhh idk :(

OpenStudy (anonymous):

this one isn't even multiple choice and we have to show work it sucks

OpenStudy (psymon):

Okay, so let me distribute everything and we'll go from there. \[25\cos(10)i + 25\sin(10)j + 30\cos(100)i + 30\sin(100)j\] Put those into the calculator and then combine the i terms and the j terms.

OpenStudy (anonymous):

so 53.3?

OpenStudy (psymon):

53.3 what? I still don't see i's and j's xD

OpenStudy (anonymous):

53.3 i hate this problem how about that? hahah

OpenStudy (psymon):

What's 25cos(10)?

OpenStudy (anonymous):

24.620

OpenStudy (psymon):

Okay, so we have 24.62i Remeber the i is there still. Okay, 25sin(10)?

OpenStudy (anonymous):

4.34j

OpenStudy (psymon):

Awesome. 30cos(100)?

OpenStudy (anonymous):

-5.2i

OpenStudy (psymon):

And finally 30sin(100) :P

OpenStudy (anonymous):

29.54j

OpenStudy (psymon):

Okay, now we combine like terms: 24.62i - 5.2i 4.34j + 29.54j

OpenStudy (anonymous):

19.41i and 33.88j

OpenStudy (psymon):

That's your answer. In the x-direction, the force felt is only 19.41 pounds, but in the y-direction 33.88 poundsof force is felt.  So this would be the overall direction of the force.